# Spatio-Temporal Graph Neural Network for Urban Spaces: Interpolating Citywide Traffic Volume

### v1

@@ 3.3 Model pipeline and loss
The neural network consists of three Diffusion Graph Convolutional Network (DGCN) layers, similar to the work of Wu et al. (2021a) and Zhuang et al. (2022) (mathematical definition of all layers and pipeline depiction in Appendix D), and an output layer. This output layer is tailored to the chosen loss function and outputs the parameters needed to compute the respective loss. In our datasets, traffic volume is characterized by a right-skewed distribution with an inflated number of zeros (Figure 1). The training loss should account for this. Mean square error (MSE) losses, commonly used in highway models, are ill-suited for this setting. They optimize for the mean and are overly sensitive to large values. This can cause the model to overestimate lower traffic volumes and underestimate higher ones. To address this, we experiment with different loss functions: we use the more robust mean absolute error (MAE) and Gaussian Negative Log-Likelihood (GNLL), as well as the Negative Binomial (NB) and Zero Inflated Negative Binomial (ZINB) losses from Zhuang et al. (2022). These are tailored to discrete, skewed data, with ZINB additionally handling excess zeros well (mathematical definition of losses in Appendix E).
@@ 3.4 Evaluation
Training, validation, and test set. We hypothesize that the amount of ground truth data, i.e. the share of street segments equipped with a sensor, affects the model’s predictive performance. To simulate different levels of data scarcity, we fix 10% of street segments for testing and vary the share of street segments available for training and validation. We start with all 90% of segments (split 9:1 for training and validation) and gradually reduce the available share to 1%. As the prediction is made at the same time as the training data, there is no temporal split.
Error metrics. We assess interpolation performance using the MAE and root mean squared error (RMSE) ( Appendix F). We avoid percentage-based metrics due to the prevalence of zeros and outliers in the data. Due to the skewed and zero-inflated nature of the data, we additionally report the Kullback-Leibler divergence (KL-divergence) to compare predicted and actual distributions, as done in related work on similar data (Zhuang et al., 2022). Since KL-divergence quantifies the disparity between two distributions, lower values are preferable. To compute the KL divergence, we require the predicted and actual probability distributions, which we approximate by discretizing the continuous values into bins. To evaluate how well the model captures the distribution’s zero-inflated nature, we compute the true-zero rate. Which measures the proportion of true zeros correctly predicted as (near-)zero values, with higher values being desirable (mathematical definition in Appendix G).
@@ 4 Data
To evaluate our model, we require urban datasets with traffic volume information at the street-segment level, combined with rich auxiliary features. However, to the best of our knowledge, no city provides complete sensor-based traffic volume data across its entire street network. Therefore, we assembled two large-scale datasets ourselves: Berlin Strava and New York City taxi data. For both, we simulate full sensor coverage by aggregating trajectory data to the street segment level, and enrich them with extensive contextual features. While we refer to sensors throughout, all evaluations are based on these proxy datasets, not actual sensor data. While Berlin data reflects bicycle traffic, NYC data captures motorized mobility, together offering a diverse and representative testbed for urban interpolation tasks.
@@ 4.1 Berlin Strava and New York taxi volume measurements
The Strava dataset contains the daily number of bike trips recorded via the Strava app (Strava Metro, 2024) at the street segment level in Berlin from 2019 to 2023. Strava aggregates publicly shared trips across sidewalks, bike paths, and vehicle lanes, rounding volumes to the nearest multiple of five for privacy. We further aggregate these measurements to the street segment level for our analysis. Although the data is biased toward young, male, and sporty users (descriptive statistics in Appendix H), it still provides valuable insights into urban cycling patterns (Kaiser et al., 2025; Lee and Sener, 2021).
The taxi dataset covers hourly taxi trips in Manhattan for January and February 2016, based on publicly available data (New York Taxi and Limousine Commission, 2016). For computational feasibility, we focus on trips within Manhattan during this period, which includes GPS coordinates for pickup and drop-off locations. From these, we reconstruct trajectories using the Open Source Routing Machine, which computes vehicle routes based on street network data, including geometry, one-way streets, and turn restrictions (Luxen and Vetter, 2011). Based on these trajectories, we compute the number of taxis passing per street segment (Boeing, 2024).
For both datasets, we remove observations deviating more than three standard deviations from a segment’s mean. Due to the strong right skew, this only removes high values. These values are caused by events like exception occasions temporarily disrupting normal traffic patterns, such as bike demonstrations or street protests. Since our focus is to model normal traffic behavior, e.g., to inform infrastructure planning, we prioritize capturing average conditions over rare anomalies. Filtering per segment preserves meaningful variation, such as consistently busy streets, while improving overall data quality. Even after filtering, the distributions remain highly right-skewed (Figure 1(d)).
@@ 4.2 Graph structure
We base the Berlin graph on the city’s official cycling network (Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin, 2024a), which includes both major traffic routes and smaller streets with high cycling volumes (Figure 1(c)). This network, smaller than Berlin’s full street network, makes the task computationally manageable and comprises 4,958 street segments. For New York, we construct the graph using all public streets in Manhattan obtained from OpenStreetMap data, resulting in a network of 8,156 street segments (Figure 1(a)).
@@ 4.3 Predictive features
We compile further features that have been helpful in previous research in predicting cycling volume and motorized traffic. We outline this related data below, and include a list of all features in Appendix I.
@@
We also include network connectivity measures, which reflect how integrated a street segment is within the overall network. Connectivity influences cycling and motorized traffic demand by shaping movement patterns through the network’s geometric properties (Hillier and Iida, 2005; Hochmair et al., 2019; Schön et al., 2024). These measures, computed based on the designated graph structure, include degree, betweenness, closeness, and clustering coefficient (mathematical definitions in Appendix J).
@@ 5.2 Different Adjacency Matrices
We find that the choice of the adjacency matrix significantly affects performance. The binary matrix consistently performs best. It achieves the lowest MAE, 7.14 for Berlin (vs. 8.92–9.73) and 23.03 for New York (vs. 30.99–40.50), as well as the lowest RMSE, 13.8 (vs. 16.8–17.4) and 56.8 (vs. 70.9–97.9), respectively. It also delivers the lowest KL divergence, 0.10 (vs. 0.12–0.21), and the highest true-zero rates: 0.603 (vs. 0.461–0.403) and 0.731 (vs. 0.516–0.006). We also test combinations of adjacency matrices, while incurring significantly higher computational costs, we find no meaningful performance gains. Full results and technical details are reported in Appendix K.
The performance advantage of binary adjacency matrices likely arises because they restrict message passing to immediate neighbors. In contrast, distance- and similarity-based matrices introduce weaker or noisier connections, diluting the relevance of the aggregated information. By focusing on clear, local dependencies, binary matrices provide a sharper and more informative signal for interpolation. Additionally, binary adjacency matrices offer clear computational advantages. Their simplicity reduces training time, making them a practical and robust choice for city-scale applications. Yet, adjacency matrices that carry traffic assignment information - and thus actual demand-network interactions - may provide improvements, but may be hard to obtain in most practical cases.
@@ 5.3 Varying data scarcity during training data
Data scarcity poses a challenge for the interpolation of urban traffic volumes. We investigate how GNNUI’s performance changes with decreasing sensor availability. This analysis uses a binary adjacency matrix. We evaluate GNNUI across sensor coverage levels from 1% to 90%—using a step size of 1 percentage point until 10%, and 10 percentage points thereafter. To reduce variability across experiments, we sample from the available ground truth sensor locations in a nested fashion: lower-coverage training sets are subsets of higher-coverage ones (e.g., the 70% set is fully contained within the 80% and 90% sets). As before, we include the entire graph during training: all node features are visible to the model, but the target variables of nodes excluded from the training set are treated as missing. That is, nodes not selected for training still appear in the graph with their features, but their true traffic volumes are masked and not used for loss computation. The test set remains fixed, using the same 10% of nodes for evaluation across all experiments. Results are shown in Figure 4. For completeness, we report results for GNNUI using distance and similarity-based adjacency matrices in Appendix L.
@@
Figure 4: Performance of GNNUI, IGNNK, and XGBoost under varying levels of data scarcity. Kindly notice that there is a structural break on the x-axis at 10%.
@@ 6 Conclusion
Our results suggest that sensor locations vary in informativeness, indicating that strategic sensor placement could enhance interpolation performance. Investigating this systematically remains an important avenue for future research. Further exploration into robust modeling techniques for urban areas with extremely limited sensor data is also warranted. We note that current benchmarking efforts for GNN-based traffic models are largely focused on highway and motorized traffic datasets. To better capture the complexity of urban environments, we advocate for systematically including diverse urban datasets and multiple transport modes in future benchmarking protocols.
@@ Acknowledgements
This section summarizes the key variables used across the appendix. Most of these are not defined in the main text but are introduced in detail in Appendix Appendix B. They are listed here for clarity and to avoid repetition.
@@
Appendix Appendix A Training Loss and Error Computation
@@
training loss masked $\text{training loss masked only}=\text{loss}\left(\hat{T}_{\text{s}},T_{\text{s}};(\mathbf{1}-M_{\text{s}})\circ N_{\text{s}}\right).$ (4)
@@
training loss masked $\text{training loss masked only}=\text{loss}\left(\hat{T}_{\text{s}},T_{\text{s}};N_{\text{s}}\right).$ (5)
@@
For the prediction of $T^{u}_{t}$, we create the corresponding matrix $W_{t}\in\mathbb{R}^{(n_{t}^{s}+n^{u}_{t})\times(n_{t}^{s}+n^{u}_{t})}$, the missing observations matrix $N_{t}\in\mathbb{R}^{(n_{t}^{s}+n^{u}_{t})\times h\times 1}$ and the masking matrix $M_{t}\in\mathbb{R}^{(n_{t}^{s}+n^{u}_{t})\times h\times 1}$. Here $M_{t}$ is equal to 0, if a node is unseen. Thus, $T_{t}=[T_{t}^{s},T^{u}_{t}]$ and $X_{t}=[X_{t}^{s},X^{u}_{t}]$
We then feed $T_{t}^{M,N}=X_{t}\circ M_{t}\circ N_{t}$ and $X_{t}=[X_{t}^{s},X^{u}_{t}]$ to the model to obtain $\hat{T}_{t}=[\hat{T_{t}^{s}},\hat{T}^{u}_{t}]$
@@
Appendix Appendix B Training Through Masking Strategy
@@
$X_{\text{s}}=X[\{i^{1},\cdots,i^{n_{0}},\cdots,i^{n_{0}+n_{m}}\},[j,j+h),:]$
$T_{\text{s}}=T[\{i^{1},\cdots,i^{n_{0}},\cdots,i^{n_{0}+n_{m}}\},[j,j+h)]$
@@
$N_{\text{s}}[i,\tilde{j}]=\begin{cases}0,&\text{if}\ \text{the measurement at }X_{\text{s}}[i,\tilde{j}]\text{ is missing,}\\ 1,&\text{otherwise.}\end{cases}$ (9)
We element-wise multiply the target variable $T_{\text{s}}$ with $N_{\text{s}}$, effectively setting missing target variables to zero. The target variable of the sample is $T^{M,N}_{\text{s}}=T_{\text{s}}\circ M_{\text{s}}\circ N_{\text{s}}$, with $\circ$ denoting element-wise multiplication. Thus, for each sample we train the model to predict $\hat{T}_{\text{s}}\in\mathbb{R}^{(n_{o}+n_{m})\times h\times 1}$ given $X_{\text{s}}\in\mathbb{R}^{(n_{o}+n_{m})\times h\times k}$ and $T^{M,N}_{\text{s}}\in\mathbb{R}^{(n_{o}+n_{m})\times h\times 1}$.
Appendix Appendix C Temporal Generalization Ablation Study

### v2

@@ 3.3 Model pipeline and loss
The neural network consists of three Diffusion Graph Convolutional Network (DGCN) layers, similar to the work of Wu et al. (2021a) and Zhuang et al. (2022) (mathematical definition of all layers and pipeline depiction in D), and an output layer. This output layer is tailored to the chosen loss function and outputs the parameters needed to compute the respective loss. In our datasets, traffic volume is characterized by a right-skewed distribution with an inflated number of zeros (Figure 1). The training loss should account for this. Mean square error (MSE) losses, commonly used in highway models, are ill-suited for this setting. They optimize for the mean and are overly sensitive to large values. This can cause the model to overestimate lower traffic volumes and underestimate higher ones. To address this, we experiment with different loss functions: we use the more robust mean absolute error (MAE) and Gaussian Negative Log-Likelihood (GNLL), as well as the Negative Binomial (NB) and Zero Inflated Negative Binomial (ZINB) losses from Zhuang et al. (2022). These are tailored to discrete, skewed data, with ZINB additionally handling excess zeros well (mathematical definition of losses in E).
@@ 3.4 Evaluation
Training, validation, and test set. We hypothesize that the amount of ground truth data, i.e. the share of street segments equipped with a sensor, affects the model’s predictive performance. To simulate different levels of data scarcity, we fix 10% of street segments for testing and vary the share of street segments available for training and validation. We start with all 90% of segments (split 9:1 for training and validation) and gradually reduce the available share to 1%. All time steps of a given segment belong to the same subset. No temporal split is used because the task is spatial interpolation rather than temporal forecasting.
Error metrics. We assess interpolation performance using the MAE and root mean squared error (RMSE) ( F). We avoid percentage-based metrics due to the prevalence of zeros and outliers in the data. Due to the skewed and zero-inflated nature of the data, we additionally report the Kullback-Leibler divergence (KL-divergence) to compare predicted and actual distributions, as done in related work on similar data (Zhuang et al., 2022). Since KL-divergence quantifies the disparity between two distributions, lower values are preferable. To compute the KL divergence, we require the predicted and actual probability distributions, which we approximate by discretizing the continuous values into bins. To evaluate how well the model captures the distribution’s zero-inflated nature, we compute the true-zero rate. Which measures the proportion of true zeros correctly predicted as (near-)zero values, with higher values being desirable (mathematical definition in G).
@@ 4 Data
To evaluate our model, we require urban datasets with traffic volume information at the street-segment level, combined with rich auxiliary features. However, to the best of our knowledge, no city provides complete sensor-based traffic volume data across its entire street network. Therefore, we assembled two large-scale datasets ourselves: Berlin Strava and New York City taxi data. For both, we simulate full sensor coverage by aggregating trajectory data to the street segment level and enriching them with extensive contextual features. While we refer to sensors throughout, all evaluations are based on these proxy datasets, not actual sensor data. While Berlin data reflects bicycle traffic, NYC data captures motorized mobility, together offering a diverse and representative testbed for urban interpolation tasks.
@@ 4.1 Berlin Strava and New York taxi volume measurements
The Strava dataset contains the daily number of bike trips recorded via the Strava app (Strava Metro, 2024) at the street segment level in Berlin from 2019 to 2023. Strava aggregates publicly shared trips across sidewalks, bike paths, and vehicle lanes, rounding volumes to the nearest multiple of five for privacy. We further aggregate these measurements to the street segment level for our analysis. Although the data is biased toward young, male, and sporty users (descriptive statistics in H), it still provides valuable insights into urban cycling patterns (Kaiser et al., 2025; Lee and Sener, 2021).
The taxi dataset covers hourly taxi trips in Manhattan for January and February 2016, based on publicly available data (New York Taxi and Limousine Commission, 2016). For computational feasibility, we focus on trips within Manhattan during this period, which includes GPS coordinates for pickup and drop-off locations. From these, we reconstruct trajectories using the Open Source Routing Machine, which computes vehicle routes based on street network data, including geometry, one-way streets, and turn restrictions (Luxen and Vetter, 2011). Based on these trajectories, we compute the number of taxis passing per street segment (Boeing, 2025).
For both datasets, we remove observations deviating more than three standard deviations from a segment’s mean. Due to the strong right skew, this only removes high values. These values are caused by events like exception occasions temporarily disrupting normal traffic patterns, such as bike demonstrations or street protests. Since our focus is to model normal traffic behavior, e.g., to inform infrastructure planning, we prioritize capturing average conditions over rare anomalies. Filtering per segment preserves meaningful variation, such as consistently busy streets, while improving overall data quality. Even after filtering, the distributions remain highly right-skewed (Figure 1).
@@ 4.2 Graph structure
We base the Berlin graph on the city’s official cycling network (Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin, 2024a), which includes both major traffic routes and smaller streets with high cycling volumes (Figure 1(b)). This network, smaller than Berlin’s full street network, makes the task computationally manageable and comprises 4,958 street segments. For New York, we construct the graph using all public streets in Manhattan obtained from OpenStreetMap data, resulting in a network of 8,156 street segments (Figure 1(a)).
@@ 4.3 Predictive features
We compile further features that have been helpful in previous research in predicting cycling volume and motorized traffic, which also contribute in our model (see J for a feature importance analysis). We outline this related data below, and include a list of all features in I.
@@
We also include network connectivity measures, which reflect how integrated a street segment is within the overall network. Connectivity influences cycling and motorized traffic demand by shaping movement patterns through the network’s geometric properties (Hillier and Iida, 2005; Hochmair et al., 2019; Schön et al., 2024). These measures, computed based on the designated graph structure, include degree, betweenness, closeness, and clustering coefficient (mathematical definitions in K).
@@ 5.1 Performance of GNNUI
In addition to the interpolation task, we evaluate GNNUI in a more challenging setting where it must predict values at new locations during future time periods. This setup allows us to assess its temporal generalisation capabilities and compare them to the literature focusing on forecasting. Although GNNUI was not specifically designed for forecasting, it still outperforms the standard baselines in this more challenging environment, demonstrating its robustness and flexibility across different tasks. See C for details.
While Strava and taxi data represent biased subsets of overall traffic, they still capture meaningful spatiotemporal patterns typical of broader mobility behavior. GNNUI’s strong performance across both datasets demonstrates its ability to learn from such patterns under real-world, noisy conditions. This suggests that GNNUI can generalize well to other types of traffic data, including more representative datasets such as general motorized traffic or pedestrian flows.
@@ 5.2 Different Adjacency Matrices
We find that the choice of the adjacency matrix significantly affects performance. The binary matrix consistently performs best. It achieves the lowest MAE, 7.14 for Berlin (vs. 8.92–9.73) and 23.03 for New York (vs. 30.99–40.50), as well as the lowest RMSE, 13.8 (vs. 16.8–17.4) and 56.8 (vs. 70.9–97.9), respectively. It also delivers the lowest KL divergence, 0.10 (vs. 0.12–0.21), and the highest true-zero rates: 0.603 (vs. 0.461–0.403) and 0.731 (vs. 0.516–0.006). We also evaluate hybrid adjacency setups, where multiple matrices are combined by training separate subnetworks for each. While this increases computational cost, it does not lead to meaningful performance gains, likely because the added edges introduce noise or less informative relationships, weakening the strong local signal captured by the binary adjacency. Full results and technical details are reported in L.
The performance advantage of binary adjacency matrices likely arises because they restrict message passing to immediate neighbors. In contrast, distance- and similarity-based matrices introduce weaker or noisier connections, diluting the relevance of the aggregated information. By focusing on clear, local dependencies, binary matrices provide a sharper and more informative signal for interpolation. Additionally, binary adjacency matrices offer clear computational advantages. Their simplicity reduces training time, making them a practical and robust choice for city-scale applications.
It should be noted that this performance advantage is demonstrated on Strava and taxi data, which represent biased subsets of total urban traffic. Nonetheless, the strong and consistent results across both datasets suggest that the relative effectiveness of different types of adjacency matrices is likely robust and transferable to other traffic domains.
@@ 5.3 Varying data scarcity during training data
Data scarcity poses a challenge for the interpolation of urban traffic volumes. We investigate how GNNUI’s performance changes with decreasing sensor availability. This analysis uses a binary adjacency matrix. We evaluate GNNUI across sensor coverage levels from 1% to 90%—using a step size of 1 percentage point until 10%, and 10 percentage points thereafter. To reduce variability across experiments, we sample from the available ground truth sensor locations in a nested fashion: lower-coverage training sets are subsets of higher-coverage ones (e.g., the 70% set is fully contained within the 80% and 90% sets). As before, we include the entire graph during training: all node features are visible to the model, but the target variables of nodes excluded from the training set are treated as missing. That is, nodes not selected for training still appear in the graph with their features, but their true traffic volumes are masked and not used for loss computation. The test set remains fixed, using the same 10% of nodes for evaluation across all experiments. Results are shown in Figure 4. For completeness, we report results for GNNUI using distance and similarity-based adjacency matrices in M. Both datasets reflect specific mobility subpopulations rather than the full spectrum of urban traffic. The consistent findings across these two distinct cases, however, suggest that the learnings are transferable to other traffic contexts.
@@
Figure 4: Performance of GNNUI, IGNNK, and XGBoost under varying levels of data scarcity. Notice that there is a structural break on the x-axis at 10%.
@@ 6 Conclusion
Our results suggest that sensor locations vary in informativeness, indicating that strategic sensor placement could enhance interpolation performance. Investigating this systematically remains an important avenue for future research. Further exploration into robust modeling techniques for urban areas with extremely limited sensor data is also warranted. We note that current benchmarking efforts for GNN-based traffic models are largely focused on highway and motorized traffic datasets. To better capture the complexity of urban environments, we advocate for systematically including diverse urban datasets and multiple transport modes in future benchmarking protocols. Lastly, future work could investigate augmenting the published Strava and taxi datasets to reduce inherent biases.
@@ Acknowledgements
This section summarizes the key variables used across the appendix. Most of these are not defined in the main text but are introduced in detail in Appendix B. They are listed here for clarity and to avoid repetition.
@@
Appendix A Training Loss and Error Computation
@@
$\text{training loss masked only}=\text{loss}\left(\hat{T}_{\text{s}},T_{\text{s}};(\mathbf{1}-M_{\text{s}})\circ N_{\text{s}}\right).$ (4)
@@
$\text{training loss masked only}=\text{loss}\left(\hat{T}_{\text{s}},T_{\text{s}};N_{\text{s}}\right).$ (5)
@@
For the prediction of $T^{u}_{t}$, we create the corresponding matrix $W_{t}\in\mathbb{R}^{(n_{t}^{s}+n^{u}_{t})\times(n_{t}^{s}+n^{u}_{t})}$, the missing observations matrix $N_{t}\in\mathbb{R}^{(n_{t}^{s}+n^{u}_{t})\times h\times 1}$ and the masking matrix $M_{t}\in\mathbb{R}^{(n_{t}^{s}+n^{u}_{t})\times h\times 1}$. Here $M_{t}$ is equal to 0, if a node is unseen. Thus, $T_{t}=[T_{t}^{s},T^{u}_{t}]$ and $X_{t}=[X_{t}^{s},X^{u}_{t}]$.
We then feed $T_{t}^{M,N}=X_{t}\circ M_{t}\circ N_{t}$ and $X_{t}=[X_{t}^{s},X^{u}_{t}]$ to the model to obtain $\hat{T}_{t}=[\hat{T_{t}^{s}},\hat{T}^{u}_{t}]$ .
@@
Appendix B Training Through Masking Strategy
@@
$X_{\text{s}}=X[\{i^{1},\cdots,i^{n_{0}},\cdots,i^{n_{0}+n_{m}}\},[t,t+h),:]$
$T_{\text{s}}=T[\{i^{1},\cdots,i^{n_{0}},\cdots,i^{n_{0}+n_{m}}\},[t,t+h)]$
@@
$N_{\text{s}}[i,\tilde{t}]=\begin{cases}0,&\text{if}\ \text{the measurement at }X_{\text{s}}[i,\tilde{t}]\text{ is missing,}\\ 1,&\text{otherwise.}\end{cases}$ (9)
We element-wise multiply the target variable $T_{\text{s}}$ with $N_{\text{s}}$, effectively setting missing target variables to zero. The target variable of the sample is $T^{M,N}_{\text{s}}=T_{\text{s}}\circ M_{\text{s}}\circ N_{\text{s}}$, with $\circ$ denoting element-wise multiplication.
To enable the model to distinguish between true zero values, masked values, and missing values, all represented numerically as zero, we explicitly provide this information as node features to the model. Specifically, the two binary matrices are included as two channels in the node feature tensor $X_{s}$. Let $X_{s}^{\text{raw}}\in\mathbb{R}^{(n_{o}+n_{m})\times h\times k^{\text{raw}}}$ denote the feature tensor without the two binary indicators. We augment it by appending $M_{s}$ (indicating masked values) and $N_{s}$ (indicating missing values) as additional channels along the feature dimension. The resulting input tensor becomes:
$X_{s}=\text{concat}(X_{s}^{\text{raw}},M_{s},N_{s})\in\mathbb{R}^{(n_{o}+n_{m})\times h\times k}$
where $k=k^{\text{raw}}+2$ reflects the inclusion of the two binary indicators. We include these binary features in the exact same manner during inference.
Thus, for each sample we train the model to predict $\hat{T}_{\text{s}}\in\mathbb{R}^{(n_{o}+n_{m})\times h\times 1}$ given $X_{\text{s}}\in\mathbb{R}^{(n_{o}+n_{m})\times h\times k}$ and $T^{M,N}_{\text{s}}\in\mathbb{R}^{(n_{o}+n_{m})\times h\times 1}$.
Appendix C Temporal Generalization Ablation Study